OpenStudy (anonymous):

How do I determine the type of quadric surface for an equation like this? x^2 + 2y^2 + 4z^2 - 4z + 1 = 0

OpenStudy (anonymous):

this is a sphere equation..

OpenStudy (anonymous):

\[(x-x_0)^2+(y-y_0)^2+(z-z_0)^2=r^2\]

OpenStudy (anonymous):

you sure about eq. because r=0

OpenStudy (anonymous):

i've figured it out, there is only 1 solution to the equation which is (0,0,1/2), which means the equation is a single point. thanks anyway!

OpenStudy (anonymous):

you mean real solution..

OpenStudy (anonymous):

it is clear..\[x^2+2y^2+(2z-1)^2=0\]